OpenStudy (anonymous):

Rosa works due north of home. Her husband Juan works due east. They leave for work at the same time. By the time Rosa is 7 miles from home, the distance between them is one mile more than Juan's distance from home. How far from home is Juan?

OpenStudy (anonymous):

Sqrt[7^2 + x^2]= x+1 7^2+x^2=(x+1)^2 49+x^2=x^2+2x+1

OpenStudy (anonymous):

2x=48 x=24

OpenStudy (anonymous):

thanks